OpenStudy (anonymous):

I have a problem with a line integral. My path is defined as an elliptic curve (x^2)/9 + (y^2)/4 ) = 1. My function is <2x(y^2) - 1,x^3> .

OpenStudy (anonymous):

Now i tried parameterising my curve, and i get the result: s = <3sin t , 2 cos t> So my ds = <dx, dy> = <3cos t, -2sin t>dt

OpenStudy (anonymous):

with\[ 0<t< \pi/2\] Now my integral looks like \[\int\limits_{0}^{\pi/2} <6sint * (2 \cos t) ²-1, (2\cos t ) ^3> * <3\cos t, -2 \sin t >dt\]

OpenStudy (anonymous):

Seems pretty odd to me... can anyone spot the mistake?

OpenStudy (anonymous):

s = <3sin t , 2 cos t> should that be s=<3Cos t, 3 Sin t> ?

OpenStudy (anonymous):

my starting point is (0, 2) so when t = 0 the x should be 0 and the y should be 2, so sin (0) = 0 and cos(0) =1 .